How many gallons of 15% alcohol solution and 45% alcohol solution must be mixed to get 12 gallons of 25% alcohol solution?
Answer by richwmiller(17219)

a+b=12,
0.15*a+0.45*b=0.25*12
a=12-b
0.15*(12-b)+0.45*b=3
1.8-0.15b+0.45*b=3
0.3*b=1.2
b=4
a=12-b
a=8 b=4
check
0.15*8+0.45*4=0.25*12
1.2+1.8=3
3=3
ok